# EditScore: Unlocking Online RL for Image Editing via High-Fidelity Reward Modeling

### v1

@@ Abstract
Instruction-guided image editing has achieved remarkable progress, yet current models still face challenges with complex instructions and often require multiple samples to produce a desired result. Reinforcement Learning (RL) offers a promising solution, but its adoption in image editing has been severely hindered by the lack of a high-fidelity, efficient reward signal. In this work, we present a comprehensive methodology to overcome this barrier, centered on the development of a state-of-the-art, specialized reward model. We first introduce EditReward-Bench, a comprehensive benchmark to systematically evaluate reward models on editing quality. Building on this benchmark, we develop EditScore, a series of reward models (7B–72B) for evaluating the quality of instruction-guided image editing. Through meticulous data curation and filtering, EditScore effectively matches the performance of learning proprietary VLMs. Furthermore, coupled with an effective self-ensemble strategy tailored for the generative nature of EditScore, our largest variant even surpasses GPT-5 in the benchmark. We then demonstrate that a high-fidelity reward model is the key to unlocking online RL for image editing. Our experiments show that, while even the largest open-source VLMs fail to provide an effective learning signal, EditScore enables efficient and robust policy optimization. Applying our framework to a strong base model, OmniGen2, results in a final model that shows a substantial and consistent performance uplift. Overall, this work provides the first systematic path from benchmarking to reward modeling to RL training in image editing, showing that a high-fidelity, domain-specialized reward model is the key to unlocking the full potential of RL in this domain. Github Link: https://github.com/VectorSpaceLab/EditScore.
@@ 1 Introduction
Despite its theoretical promise, applying online RL to high-resolution image editing remains a formidable and largely unsolved challenge (Wei et al., 2025; Wu et al., 2025a; Ahmadi et al., 2025). The primary obstacle lies in the absence of a suitable reward function: a reliable, efficient, and scalable oracle that can accurately score the quality of an edit given an instruction. A natural consideration would be to employ state-of-the-art, general-purpose Visual Language Models (VLMs) like GPT-5 (OpenAI, 2025) or other proprietary APIs as reward providers. However, this approach is impractical for online RL, with these large VLMs are expensive to query at scale. The second alternative, using powerful open-source VLMs, resolves cost issues. However, our investigations reveal a critical performance gap. As we will demonstrate, even Qwen2.5-VL-72B (Bai et al., 2025) fails to provide a sufficiently accurate and consistent reward signal to effectively guide the policy. Directly using such models as reward functions often causes unstable training or policy collapse, showing that scale alone cannot replace specialized accuracy. Progress thus hinges on an accurate reward model tailored to image editing, yet the community still lacks a powerful open-source option—posing a major barrier to advancing RL-based editing.
@@ 3.3 Annotation Pipeline
To ensure the reliability of EditReward-Bench, we designed a rigorous human annotation pipeline. The process was conducted exclusively by experts in generative AI to guarantee data quality and consistency, as illustrated in Figure 1.
Before the formal annotation, all raters scored a small subset of the entire sample pool to discuss disagreements and collectively established unified scoring guidelines. Then the raters proceeded to the formal annotation of the remaining samples. For each input with five output images randomly sampled from the candidate pool, they were asked to rank them. Our protocol uniquely allowed them to group outputs of similar quality into the same tiers; for instance, a ranking of $3|12|45$ places output 3 in the top tier, outputs 1 and 2 tied in the second, and 4 and 5 tied in the third. Crucially, each data point was annotated by two raters, and only entries where their rankings were in complete agreement were accepted into the final benchmark, ensuring maximum reliability.
These tiered rankings were then decomposed into pairwise preference tuples. Following the $3|12|45$ example, this conversion yields preference pairs where a higher-tiered item is preferred over a lower-tiered one, such as $(3,1)$, $(3,2)$, $(1,4)$ and $(2,5)$. This method correctly excludes comparisons between items within the same tier (e.g., $(1,2)$), as they represent equal quality. This process resulted in a large-scale, high-fidelity dataset of pairwise comparisons, comprising 944 pairs for prompt following, 890 for consistency, and 1,238 for overall quality, for a total of 3,072 preference pairs. A detailed overview of our custom-built annotation interface is provided in Appendix F.
@@ 4.1.1 The Approach to Reward Modeling
Inspired by (Wang et al., 2025b; a; Wu et al., 2025c), we formulate reward modeling as a conditional textual generation task. We fine-tune models from the powerful Qwen2.5-VL series (Bai et al., 2025) on a standard autoregressive objective to act as specialized evaluators. Our model EditScore takes (Instruction, Input Image, Output Image) as input, (Reasoning, Scalar Score) as output. The chain-of-thought (Wei et al., 2022) process of output not only enhances the model’s interpretability but also demonstrably improves the accuracy of the final scores.
@@ 4.1.2 Inference-time ensembling Strategy
Recent work (Snell et al., 2024; Liu et al., 2025c) has shown that the performance of LLMs can be substantially improved by increasing the computational budget at test time. Building on this insight, we introduce an inference-time ensembling strategy for our generative reward model to systematically boost its evaluation accuracy and enhance reward stability. Our approach is straightforward: for a given input triplet $\mathbf{z}=(\text{Instruction, Input Image, Output Image})$, we perform $K$ independent, stochastic forward passes through EditScore, and each pass $i$ generates a pair $(\text{reasoning}_{i},s_{i})$. Then we aggregate only the scalar scores $\{s_{1},s_{2},\dots,s_{K}\}$ to compute the final ensembled score, $S_{final}$. An intuitive explanation (Liu et al., 2025c) for this effectiveness is that each of the $K$ generated reasonings can be viewed as distinct judgment perspectives. Aggregating the scores derived from these diverse perspectives allows the final reward to more accurately reflect the true quality of an edit, leading to significant scaling effectiveness.
@@ 4.2 Data Construction Pipeline
Step I: Images selection and instructions creation. We start by selecting and filtering a diverse set of high-quality images as editing inputs. Next, we construct a series of reference instructions for various editing tasks. Using the input images and random selected reference instructions as guidance, we prompt Qwen-2.5-VL-72B to generate editing instructions that are task-consistent with the references. After generating a large set of image-instruction pairs, we applying a K-center greedy algorithm (Sener & Savarese, 2018) to select 1000 semantically diverse samples per task for reward generation.
@@ 5.1 Experimental Setup
Our final EditScore model is obtained by fine-tuning Qwen2.5-VL with LoRA (Hu et al., 2022) on our curated dataset (see Section 4.2). For evaluation, we adopt the VIEScore (Ku et al., 2024) prompt template with two key modifications: (i) enforcing a reasoning-before-scoring format, (ii) expanding the score range from $[0,10]$ to $[0,25]$. Following the formulation of VIEScore, we derive Prompt Following ($S_{PF}$) and Consistency ($S_{C}$) from its Semantic Consistency metric $S_{SC}$, while Overall Quality ($O$) is directly taken from the final score $S_{final}$. The detailed prompt templates are shown in Appendix H.
Table 2: Benchmark results on EditReward-Bench, reporting both overall pairwise accuracy and a fine-grained breakdown across four categories of edit capabilities. Notably, EditScore achieves superior performance even with its compact 7B size. Avg@4 denotes the average score over 4 forward passes.
@@ 5.3 Effective Inference-time scaling of EditScore
Figure 2: Self-ensembling offers a superior efficiency-performance trade-off compared to simply scaling model parameters. The colored solid lines show the performance scaling of our 7B, 32B, and 72B models as the number of ensemble passes (K) increases. The gray dashed line connects the single-pass (K=1) performance of these models, serving as a baseline for scaling model size alone. The results clearly indicate that scaling the number of forward passes yields a significantly higher accuracy gain per unit of computational cost.
As shown in Table 2, both scaling model size and inference-time compute improve the performance of EditScore. To further investigate which dimension contributes more effectively, we normalize the comparison by using FLOPs as a proxy for inference cost. Figure 2 presents the scaling trends of EditScore across different model sizes: colored solid lines denote inference-time compute scaling, while gray dotted lines indicate parameter scaling. We observe that, across all model sizes, scaling inference-time compute yields greater performance gains than scaling parameters. Moreover, in practical deployment, the self-ensemble strategy allows the profiling stage to be shared within the ensemble group, which improving efficiency further.
@@ 6.1 Experimental Setup
RL Training. OmniGen2 Wu et al. (2025b) is chosen as the base model due to its strong potential for further improvement, as evidenced in the best-of-$N$ experiments (Section 6.2). We employ Flow-GRPO Liu et al. (2025a) with the following hyperparameters: sampling steps $T=20$, group size $G=12$, number of unique prompts = 24, noise level $\sigma=0.9$, and KL weight $\beta=0.04$. We reformulate the standard Flow-GRPO equation to align with OmniGen2’s definition; detailed derivations are provided in Appendix B and C.
6.2 Validating EditScore Utility via best-of-$N$ Section
@@
6.3 Main RL Results and Analysis
Having established the utility of EditScore in a controlled best-of-$N$ setting and identified OmniGen2 as a high-potential candidate, we now proceed to the core of our investigation: leveraging EditScore as a direct reward signal for online RL fine-tuning.
@@ 6.3.1 Unlocking Stable RL with a High-Fidelity Reward Signal
Our experiments reveal that successful RL fine-tuning in the image editing domain is not a given; it hinges critically on the quality of the reward signal.
The Decisive Role of Reward Model Fidelity. The central thesis of our work is that a high-fidelity, specialized reward model is a prerequisite for effective RL. Figure 3(b) provides a stark validation of this claim. When attempting to use general-purpose, off-the-shelf VLMs (Qwen2.5-VL series) as reward providers, the training process is highly unstable and the policy model fails to converge, even with the 72B parameter model. This failure highlights a crucial insight: raw VLM capabilities do not readily translate into a usable reward signal. In stark contrast, our distilled EditScore provides a clean, consistent, and informative reward, enabling a stable and highly effective learning trajectory.
Unlocking Further Gains with Self-Ensembling. Beyond the model itself, the inference-time strategy for deploying the reward model is also impactful. Figure 3(c) compares the learning curves using a single-pass EditScore versus our 4-pass self-ensemble strategy. The ensembled reward signal demonstrably leads to faster convergence and a superior final policy performance. This confirms a vital connection: improvements in the reward model’s standalone accuracy (as shown in our main results) directly translate to more efficient and effective downstream RL training.
@@ 6.4 Analysis of Annotation Sources and Reward Variance
6.4.1 Quantifying the Final Performance Gains
Table 5: Performance comparison of models with RL enhancement, where our EditScore provides substantial gains for OmniGen2 across two widely adopted benchmarks.
Model GEdit-Bench-EN ImgEdit-Bench
SC PQ O O
UniPic2-SD3.5M-Kontext 6.31 3.95
UniPic2-SD3.5M-Kontext RL w/ GPT-4.1 6.55 (+0.24) 4.04 (+0.09 )
OmniGen2 6.72 7.20 6.28 3.40
OmniGen2 RL w/ EditScore 7.20 (+0.48) 7.46 (+0.26) 6.68 (+0.40) 3.63 (+0.23)
Table 5 summarizes the substantial gains achieved by OmniGen2 after RL fine-tuning with EditScore. We also provide a visualization of the results before and after RL in the Appendix 5.
On GEdit-Bench-EN, RL raises the overall score from 6.28 to 6.68, a notable improvement of +0.40, with particularly strong gains in Semantic Control (+0.48). On the more challenging ImgEdit-Bench, we also observe a meaningful increase of +0.23. For comparison, we include UniPic2 with EditReward, showing that the improvements achieved by EditScore are both competitive and substantial. These results provide strong evidence that a high-fidelity reward model such as EditScore is essential for unlocking the full potential of online RL in image editing.

### v3

@@ Abstract
Instruction-guided image editing has achieved remarkable progress, yet current models still face challenges with complex instructions and often require multiple samples to produce a desired result. Reinforcement Learning (RL) offers a promising solution, but its adoption in image editing has been severely hindered by the lack of a high-fidelity, efficient reward signal. In this work, we present a comprehensive methodology to overcome this barrier, centered on the development of a state-of-the-art, specialized reward model. We first introduce EditReward-Bench, a comprehensive benchmark to systematically evaluate reward models on editing quality. Building on this benchmark, we develop EditScore, a series of reward models (7B–72B) for evaluating the quality of instruction-guided image editing. Through meticulous data curation and filtering, EditScore effectively matches the performance of learning proprietary VLMs. Furthermore, coupled with an effective self-ensemble strategy tailored for the generative nature of EditScore, our largest variant even surpasses GPT-5 in the benchmark. We then demonstrate that a high-fidelity reward model is the key to unlocking online RL for image editing. Our experiments show that, while even the largest open-source VLMs fail to provide an effective learning signal, EditScore enables efficient and robust policy optimization. Applying our framework to a strong base model, OmniGen2, results in a final model that shows a substantial and consistent performance uplift. Overall, this work provides the first systematic path from benchmarking to reward modeling to RL training in image editing, showing that a high-fidelity, domain-specialized reward model is the key to unlocking the full potential of RL in this domain. Our code, models, data and benchmark will be released publicly.
@@ 1 Introduction
Despite its theoretical promise, applying online RL to high-resolution image editing remains a formidable and largely unsolved challenge (Wei et al., 2025; Wu et al., 2025a; Ahmadi et al., 2025). The primary obstacle lies in the absence of a suitable reward function: a reliable, efficient, and scalable oracle that can accurately score the quality of an edit given an instruction. A natural consideration would be to employ state-of-the-art, general-purpose Visual Language Models (VLMs) like GPT-5 (OpenAI, 2025) or other proprietary APIs as reward providers. However, this approach is impractical for online RL, with these large VLMs are expensive to query at scale. The second alternative, using powerful open-source VLMs, resolves cost issues. However, our investigations reveal a critical performance gap. As we will demonstrate, even Qwen2.5-VL-72B (Bai et al., 2025b) fails to provide a sufficiently accurate and consistent reward signal to effectively guide the policy. Directly using such models as reward functions often causes unstable training or policy collapse, showing that scale alone cannot replace specialized accuracy. Progress thus hinges on an accurate reward model tailored to image editing, yet the community still lacks a powerful open-source option—posing a major barrier to advancing RL-based editing.
@@ 3.3 Annotation Pipeline
To ensure the reliability of EditReward-Bench, we designed a rigorous human annotation pipeline conducted exclusively by experts in generative AI, as illustrated in Figure 1.
To ensure high-quality ground truth, we implemented a Two-Annotator Discussion Protocol. Unlike standard crowdsourcing where raters work in isolation, our protocol assigns two experts to each sample. They engage in real-time discussion to analyze visual artifacts (e.g., attribute drift) and resolve discrepancies, finalizing the ranking only upon reaching a joint consensus. We validated this protocol through a controlled study comparing it against independent labeling. The results demonstrate that the discussion-based approach significantly reduces annotation noise: for instance, in the Consistency dimension, the inter-annotator convergence rate (at a 100% agreement threshold) improved by 12.12%. Furthermore, the rankings determined by a single annotator pair achieved over 97% consistency with the majority vote derived from the full expert pool, confirming that our efficient two-annotator setup yields results highly aligned with collective expert judgment. A detailed disagreement analysis is provided in Appendix B.
Following this protocol, for each input with five output images randomly sampled from the candidate pool, the annotators were asked to rank them. Our system uniquely allowed them to group outputs of similar quality into the same tiers; for instance, a ranking of $3|12|45$ places output 3 in the top tier, outputs 1 and 2 tied in the second, and 4 and 5 tied in the third.
These tiered rankings were then decomposed into pairwise preference tuples. Following the $3|12|45$ example, this conversion yields preference pairs where a higher-tiered item is preferred over a lower-tiered one, such as $(3,1)$, $(3,2)$, $(1,4)$ and $(2,5)$. Comparisons between items within the same tier (e.g., $(1,2)$) are excluded as they represent equal quality. This process resulted in a large-scale, high-fidelity dataset of 3,072 preference pairs, comprising 944 for prompt following, 890 for consistency, and 1,238 for overall quality. For the evaluation of reward models, we utilize the preference prediction accuracy metric, calculated by the proportion of pairs $(A,B)$ where the model scores the human-preferred output higher, i.e., $S(A)>S(B)$.
@@ 4.1.1 The Approach to Reward Modeling
Inspired by (Wang et al., 2025b; a; Wu et al., 2025c), we formulate reward modeling as a conditional textual generation task. We fine-tune models from the powerful Qwen2.5-VL series (Bai et al., 2025b) on a standard autoregressive objective to act as specialized evaluators. Our model EditScore takes (Instruction, Input Image, Output Image) as input, (Reasoning, Scalar Score) as output. The chain-of-thought (Wei et al., 2022) process of output not only enhances the model’s interpretability but also demonstrably improves the accuracy of the final scores.
@@ 4.1.2 Inference-time ensembling Strategy
Recent work (Snell et al., 2024; Liu et al., 2025c) has shown that the performance of LLMs can be substantially improved by increasing the computational budget at test time. Building on this insight, we introduce an inference-time ensembling strategy for our generative reward model to systematically boost its evaluation accuracy and enhance reward stability. Our approach is straightforward: for a given input triplet $\mathbf{z}=(\text{Instruction, Input Image, Output Image})$, we perform $K$ independent, stochastic forward passes through EditScore, and each pass $i$ generates a pair $(\text{reasoning}_{i},s_{i})$. Then we aggregate only the scalar scores $\{s_{1},s_{2},\dots,s_{K}\}$ to compute the final ensembled score, $S_{final}$, using the arithmetic mean:
$S_{final}(\mathbf{z})=\frac{1}{K}\sum_{i=1}^{K}s_{i},\quad\text{where }(\text{reasoning}_{i},s_{i})=\text{EditScore}(\mathbf{z}).$ (1)
An intuitive explanation (Liu et al., 2025c) for this effectiveness is that each of the $K$ generated reasonings can be viewed as distinct judgment perspectives. Aggregating the scores derived from these diverse perspectives allows the final reward to more accurately reflect the true quality of an edit, leading to significant scaling effectiveness.
@@ 4.2 Data Construction Pipeline
Step I: Images selection and instructions creation. We start by selecting and filtering a diverse set of high-quality images as editing inputs. Next, we construct a series of reference instructions for various editing tasks. Using the input images and random selected reference instructions as guidance, we prompt Qwen-2.5-VL-72B to generate editing instructions that are task-consistent with the references. After generating a large set of image-instruction pairs, we applying a K-center greedy algorithm (Sener and Savarese, 2018) to select 1000 semantically diverse samples per task for reward generation.
@@ 5.1 Experimental Setup
Our final EditScore model is obtained by fine-tuning Qwen2.5-VL with LoRA (Hu et al., 2022) on our curated dataset (see Section 4.2). For evaluation, we adopt the VIEScore (Ku et al., 2024) prompt template with two key modifications: (i) enforcing a reasoning-before-scoring format, (ii) expanding the score range from $[0,10]$ to $[0,25]$. Following the formulation of VIEScore, we derive Prompt Following ($S_{PF}$) and Consistency ($S_{C}$) from its Semantic Consistency metric $S_{SC}$, while Overall Quality ($O$) is directly taken from the final score $S_{final}$. The detailed prompt templates are shown in Appendix M.
Table 2: Benchmark results on EditReward-Bench, reporting both overall pairwise accuracy and a fine-grained breakdown across four categories of edit capabilities. Pairwise accuracy measures the proportion of pairs where the model correctly assigns a higher reward score to the human-preferred output. Notably, EditScore achieves superior performance even with its compact 7B size. Avg@4 denotes the average score over 4 forward passes.
@@ 5.3 Effective Inference-time scaling of EditScore
(a) TFLOPs vs. Accuracy
(b) Latency vs. Accuracy
Figure 2: Self-ensembling offers a superior efficiency-performance trade-off compared to simply scaling model parameters. (a) TFLOPs vs. Accuracy: The colored solid lines (increasing $K$) show a steeper performance trajectory than the gray dashed line (scaling model size), indicating a higher return on computational investment. (b) Latency vs. Accuracy: Real-world measurements confirm that the latency cost of self-ensembling is sublinear due to shared KV-cache prefilling. Notably, EditScore-7B ($K=4$) achieves higher accuracy than the single-pass EditScore-32B ($K=1$) while requiring significantly less latency and hardware resources.
As shown in Table 2, both scaling model size and inference-time compute improve the performance of EditScore. To further investigate which dimension contributes more effectively, we normalize the comparison by using FLOPs as a proxy for inference cost. Figure 2 analyzes the scaling properties of EditScore from two perspectives: theoretical computational cost (TFLOPs, Left) and real-world inference latency (Right). In both plots, colored solid lines denote inference-time compute scaling (increasing ensemble size $K$), while the gray dashed lines represent parameter scaling (increasing model size from 7B to 72B). We observe that scaling inference-time compute yields significantly greater marginal gains than scaling model parameters. Moreover, as shown in Figure 2(b), the wall-clock latency grows sublinearly with $K$. This efficiency stems from the shared prefill stage across ensemble members within the attention mechanism, minimizing redundant computation. For a comprehensive breakdown of memory consumption, system throughput, and normalized hardware budgets, please refer to Appendix C.
@@ 6.1 Experimental Setup
RL Training. Our main experiments are conducted using OmniGen2 Wu et al. (2025b), while we also validate the generalization of our method on Flux-Kontext-dev Batifol et al. (2025). Flow-GRPO Liu et al. (2025a) is employed with hyperparameters: sampling steps $T=20$, group size $G=12$, number of unique prompts =24=24, noise level $\sigma=0.9$, and KL weight $\beta=0.04$. Ablation studies on hyperparameters are provided in Appendix E. We also reformulate the standard Flow-GRPO equation to align with OmniGen2’s notation; detailed derivations are given in Appendices F and G.
6.2 Validating EditScore Utility via best-of-NNection
@@
6.3 Online Reinforcement Learning with EditScore
We now turn to the core application of our framework: using EditScore as a direct reward signal for online RL fine-tuning. This section systematically isolates the impact of reward quality, examines training stability, and evaluates generality across models and algorithms. We additionally provide qualitative visual comparisons in Appendix I.
@@ 6.3.1 Unlocking Stable RL with a High-Fidelity Reward Signal
To isolate the contribution of the reward signal, we conducted a controlled study using OmniGen2 as the policy model while varying the source of supervision. As summarized in Table 4, raw instruction-tuned VLMs prove ineffective as verifiers; even the large-scale Qwen2.5-VL-72B fails to provide meaningful guidance, resulting in unstable optimization trajectories (see Figure 3(b)). This underscores that parameter scale alone cannot substitute for task-specific alignment. In contrast, our specialized EditScore-7B not only stabilizes training but also demonstrates competitiveness with the proprietary GPT-4.1, surpassing it on the out-of-distribution EmuEdit benchmark (e.g., 0.900 vs. 0.884 CLIP-I). Furthermore, our approach benefits directly from stronger backbones: EditScore-8B (adapted from Qwen3-VL-8B (Bai et al., 2025a)) achieves the highest overall performance across metrics, confirming the method’s scalability with advancing VLMs.
Beyond the reward model, we further investigate the impact of test-time compute. Figure 3(c) compares policy learning curves guided by a single-pass EditScore versus a 4-pass self-ensemble strategy (Avg@4). The ensembled signal significantly improves evaluation accuracy, leading to faster convergence and superior final performance. As we discussed in Appendix L, single pass evaluation occasionally produces unstable reward signal, while self-ensemble alleviate it, thus providing a consistent gradient for the policy to traverse.
Table 4: Controlled Evaluation of Reward Models. Using OmniGen2 as the fixed base, we compare policies trained with different reward signals. EditScore variants consistently outperform both raw VLMs and GPT-4.1, even on the out-of-distribution (OOD) EmuEdit benchmark. Note that raw Qwen2.5-VL baselines fail to provide effective guidance.
Reward Signal GEdit-Bench ImgEdit EmuEdit (OOD)
SC PQ Overall Overall CLIP-I CLIP-O DINO
Baseline (No RL) 6.72 7.20 6.28 3.40 0.876 0.309 0.822
RL w/ Qwen2.5-VL-7B 6.87 6.86 6.22 - - - -
RL w/ Qwen2.5-VL-72B 6.89 7.21 6.42 3.60 0.871 0.311 0.814
RL w/ GPT-4.1 7.24 7.41 6.73 3.66 0.884 0.312 0.843
RL w/ EditScore-7B (Avg@4) 7.20 7.46 6.68 3.63 0.900 0.310 0.870
RL w/ EditScore-Qwen3-VL-8B (Avg@4) 7.27 7.77 6.89 3.62 0.891 0.310 0.861
6.4 Generality Across Editors and Algorithms
Table 5: Generality Analysis. Left: Consistent improvements on a different backbone (Flux-Kontext-dev). Right: Compatibility with advanced RL algorithms (TempFlow-GRPO (He et al., 2025)).
(a) Flux-Kontext-dev Backbone
Method SC PQ O
Baseline 6.59 7.61 6.15
RL w/ Ours 7.16 7.95 6.87
(b) Algorithm Comparison (OmniGen2 base)
Algorithm SC PQ O
Flow-GRPO 7.20 7.46 6.68
TempFlow-GRPO 7.53 7.99 7.21
We further assess the robustness of EditScore by varying both the underlying generative backbone and the RL optimization algorithm. Throughout the following experiments, we use EditScore-7B (Avg@4) as the reward signal.
Generality Across Editors. A reliable reward model should capture intrinsic editing quality rather than overfit to the behavior of a particular policy. To test this, we replace OmniGen2 with Flux-Kontext-dev (Batifol et al., 2025), a substantially different editing model. As shown in Table 5 (Left), EditScore-based RL consistently yields strong improvements across all metrics (e.g., +0.57 in SC and +0.72 in Overall). These results indicate that EditScore provides a model-agnostic, universally useful reward signal that effectively enhances editing policies regardless of their initialization.
Compatibility with Advanced RL Algorithms. We also study whether the reward landscape defined by EditScore supports more advanced optimization strategies. Beyond the standard Flow-GRPO, we evaluate TempFlow-GRPO (He et al., 2025), which introduces temporally aware loss reweighting to emphasize low–signal-to-noise segments during training. As shown in Table 5 (Right), combining EditScore with TempFlow-GRPO delivers further gains, achieving an Overall score of 7.21. This demonstrates that our reward signal is sufficiently stable and high-fidelity to benefit from stronger search algorithms, scaling gracefully without causing reward hacking or optimization instability.